(2y^3-x^3)dx+3x^2ydy=0
1
Expert's answer
2021-07-12T13:24:23-0400

Given "(2y^3-x^3)dx+3x^2ydy=0"

"\frac{dy}{dx} = -\frac{(2y^3-x^3)}{3x^2y} = \frac{(x^3-2y^3)}{3x^2y}"


Let y = vx, then "\frac{dy}{dx } = v+x\frac{dv}{dx}"


"v+x\frac{dv}{dx} = \frac{(x^3-2v^3x^3)}{3x^3v} = \frac{1-2v^3}{3v}"


"x\frac{dv}{dx} = \frac{1-2v^3}{3v}-v = \frac{1-2v^3-3v^2}{3v}"


"\frac{3vdv}{1-2v^3-3v^2} = \frac{dx}{x}"



Integrating both sides,

"\int \frac{3vdv}{1-2v^3-3v^2} =\int \frac{dx}{x}"


Doing partial fractions,

"\frac{3v}{1-2v^3-3v^2} = \frac{1}{9(v+1)} -\frac{1}{3(v+1)^2}-\frac{2}{9(2v-1)}"


Now integrate both sides,

"\int \frac{1}{9(v+1)} -\frac{1}{3(v+1)^2}-\frac{2}{9(2v-1)} dv =\int\frac{dx}{x}"


"\frac{1}{3}ln|v+1| +\frac{1}{v+1} -\frac{1}{3}ln|2v-1| = ln|x| + ln|C|"


"\frac{1}{3}ln|\frac{v+1}{2v-1}| +\frac{1}{v+1} = ln|Cx|"


Simply putting "v = \frac{y}{x}"


"\frac{1}{3}ln|\frac{y+x}{2y-x}| +\frac{x}{y+x} = ln|Cx|"